OpenStudy (anonymous):

4(x − 1) − (x + 2) = 12

OpenStudy (lilsis76):

you would mult. the 4 and - into the parenthesis 4x-4-x-2=12 combine like terms 3x-6=12 3x=6 divide by 3 x=2

OpenStudy (anonymous):

x = 6. Distribute the 4 to (x-1) and distribute the - to (x+2), combine like terms to get 3x-6=12, subtract 6 from each side, divide by 3, and you're left with 6. Plug 6 back into the problem to check your answer.